# E-Calib: A Fast, Robust and Accurate Calibration Toolbox for Event Cameras

### v1

@@ Abstract
Event cameras triggered a paradigm shift in the computer vision community delineated by their asynchronous nature, low latency, and high dynamic range. Calibration of event cameras is always essential to account for the sensor intrinsic parameters and for 3D perception. However, conventional image-based calibration techniques are not applicable due to the asynchronous, binary output of the sensor. The current standard for calibrating event cameras relies on either blinking patterns or event-based image reconstruction algorithms. These approaches are difficult to deploy in factory settings and are affected by noise and artifacts degrading the calibration performance. To bridge these limitations, we present E-Calib, a novel, fast, robust, and accurate calibration toolbox for event cameras utilizing the asymmetric circle grid, for its robustness to out-of-focus scenes. The proposed method is tested in a variety of rigorous experiments for different event camera models, on circle grids with different geometric properties, and under challenging illumination conditions. The results show that our approach outperforms the state-of-the-art in detection success rate, reprojection error, and estimation accuracy of extrinsic parameters.
@@ I Introduction
Over the past decades, imaging sensors have evolved at a rapid pace providing solutions for various intelligent perception algorithms. The event camera is a bio-inspired sensor whose pixels are smart and independent of each other: when a pixel detects a change of intensity (usually triggered by motion), that pixel triggers an event, which records the coordinates of the pixel spiking, the time at which the event occurred, and the sign of the intensity change. The output of an event camera is therefore not frames at constant time intervals but rather an asynchronous stream of events in space and time. Due its asynchronous and differential nature, an event camera has three key advantages for robot manufacturing: sub-millisecond latency, microsecond resolution, and very high dynamic range [1, 2, 3]. These features grant the sensor robustness against motion blur and poorly lighted environments, marking a new paradigm for vision sensors. With such unprecedented capabilities in hand, event-based vision unveiled its impact in space robotics [4, 5, 6], robotic manufacturing [7], fast object detection and tracking [8, 9, 10, 11], autonomous navigation [12, 13, 14], and indoor positioning systems [15].
@@
More sophisticated event camera models, also called Dynamic Active-Pixel Vision Sensor (DAVIS) [18], are accompanied by frame-based imaging capabilities, where the active-pixel sensor (APS) frames are utilized for calibration using conventional calibration techniques. As event-based vision is gradually gaining maturity, recent models of event cameras are rather manufactured as stand-alone sensors. Thus, Huang et al. [19] proposed a multi-segment optimization-based approach to obtain the event camera intrinsics from the asynchronous events using the asymmetric circle grid for its robustness against defocused scenes. However, the approach utilizes the event stream as control points for the calibration optimization. Hence, sub-pixel localization accuracy of the calibration feature points is not attained degrading the calibration accuracy. This ultimately instigates a surging demand for a robust, event-driven calibration tool to obtain the intrinsic parameters of event cameras.
@@
    We validated the proposed work in rigorous experiments for different event camera models using different calibration patterns and under challenging lighting conditions. The results show that our approach outperforms the state-of-the-art in terms of detection success rate, reprojection error, extrinsics estimation accuracy, and computational efficiency.
@@ I-A Related Work
Although image processing techniques have advanced in the past decades, the unconventional nature of event cameras inhibits the applicability of such algorithms for event-based vision. Still, knowledge of the sensor intrinsic parameters remains necessary for developing robust event-based perception algorithms. The first method developed for calibrating event cameras utilized blinking calibration patterns [16]. Since event-based vision sensors require light intensity variation to generate events, blinking screens alleviate this restriction by consistently firing events resembling checkerboard and circle calibration grids. Nonetheless, such approach needs unnecessary instrumentation as specialized screens are required to render the calibration pattern at specific frequencies [17]. More importantly, it is usually desired to move the calibration pattern instead of the event camera during calibration. This is the case when the imaging sensor is mounted on unmanned aerial and ground vehicles [29], where manipulating the blinking screen is not feasible. Due to the limitations of blinking calibration patterns, recent advances in deep learning have been utilized to reconstruct grayscale frames, where conventional calibration methods are utilized to obtain the event camera intrinsics from the reconstructed frames. The first deep neural network to provide such utility is the E2VID network [30] where E2Calib [17] was proposed utilizing E2VID for the purpose of sensor calibration. Consequently, several improved networks have been developed: 1- Spade-E2VID by Cadena et al. [31] and 2- FireNet by Scheerlinck et al. [32]. However, the reconstructed images can suffer from artifacts as the sensor sweeps the calibration pattern degrading the feature extraction and calibration performance. This also shows that traditional image-based perception algorithms do not maximize the potential of event-based vision. Since traditional image processing techniques for event camera calibration do not provide the sufficient calibration accuracy, Huang et al. [19] proposed a multi-segment based optimization to obtain the sensor intrinsics directly from the asynchronous events using the asymmetric circle grid for its robustness against defocused scenes [33]. The aforementioned method involves utilizing density-based spatial clustering with applications of noise (DBSCAN) [34] for associating the events to their corresponding circles and each of the clustered events are utilized as control points in the multi-segment optimization framework. While the approach is the first method to propose an event-based framework for event camera calibration, it is limited by a couple of drawbacks. First, DBSCAN is performed on the spatial stamps of events and ignores their temporal dimension, leading to sub-optimal clustering performance. Second, the method utilizes the asynchronous events as the control points for the calibration optimization instead of the image centers of the circle targets. Hence, sub-pixel localization accuracy of the control points is not achieved with the edge events, which is directly correlated with degraded calibration performance. Finally, the method assumes that the clustered events corresponding to the circle targets are free from noise. However, this assumption is not valid as event-based sensors are characterized with high levels of measurement noise leading to sub-optimal calibration optimization.
Figure 1: A high-level block diagram of the proposed calibration method. The feature points of the pattern circles are extracted by means of efficient weighted least squares and time synchronization. The extracted centers are fed to the calibration optimization block to obtain the sensor intrinsics.
@@ II Framework Overview
An illustrative block diagram of our proposed calibration approach is shown in Fig. 1, involving a moving event camera observing an asymmetric circle grid. We chose the asymmetric circle grid as the calibration pattern for two reasons. First, the asymmetric circles calibration pattern is robust against out of focus scenes [33]. Second, the features generated from the circles are motion invariant, unlike checkerboards and AprilTag Markers where edges parallel to camera motion do not fire events [35], see Fig. 2. As the event camera sweeps the calibration pattern of known $M$ circles, the fired events need to be associated to their corresponding circles. Thus, we devise spatiotemporal density-based clustering with applications to noise (ST-DBSCAN) [36] to cluster the asynchronous events to their perspective circular features, outlined in section III-B. Unlike DBSCAN [34] which ignores the timestamps of the events, ST-DBSCAN performs optimal event-based clustering due to its capability to operate on spatiotemporal data and adhere to the sparse output of the event camera. It is worth mentioning that ST-DBSCAN requires predefined parameters $\epsilon_{s}$ and $\epsilon_{t}$, defined as the neighborhood search radius, which require tuning for different event camera models. To circumvent this limitation, the asynchronous events are normalized with respect to the sensor resolution prior to ST-DBSCAN. This is essential to perform clustering in a unified spatiotemporal domain and ensure consistent clustering performance even for different sensor sizes.
@@
After successfully associating the events to their perspective circles, the image centers of the calibration targets need to be extracted with sub-pixel localization accuracy to optimally calibrate the sensor. Such feature is not obtained by utilizing the raw events as control points for calibration optimization. A naive approach is to rely on ST-DBSCAN cluster centers as the feature points of the calibration targets. However, due to the inherent noisy nature of the event camera, the cluster centers drift away from the actual circle centers. Even though ST-DBSCAN is characterized with robustness to noise, noisy events can still be mapped to the circle targets, especially when they are close to the real events. On the other hand, the observed circular features form dense slanted cylinders in spatiotemporal space due to monotonically increasing events temporal stamps, preventing the applicability of image-based feature extraction methods. To extract the image points of the pattern circles, each clustered set of events is first fed to the efficient weighted least squares (eRWLS) algorithm, discussed in section IV-A, where cylinders are robustly fitted to the events clusters. It is essential to highlight that eRWLS assumes the probability density function (PDF) of the residual vector is dominated by a normal distribution. Accordingly, eRWLS efficiently reweights the residual vector while penalizing the outlier weights iteratively until convergence. Finally, since background clutter can be observed, a modified hierarchical clustering algorithm, detailed in section IV-B, is devised to extract the calibration pattern in the image plane apart from the background given the geometric properties of the asymmetric circle grid. As soon as the image centers are identified, they are fed to the calibration optimizer along with the pattern geometric properties to obtain the sensor intrinsic parameters.
@@ III-A Event-Based Vision
$s\begin{bmatrix}u\\ v\\ 1\end{bmatrix}=K\begin{bmatrix}R\text{ }T\end{bmatrix}\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix},$ (2)
where $x$, $y$, $z$ represent the world coordinates of a point in 3D space mapped to image coordinates $u$ and $v$. $[R\text{ }T]$ is the transformation from world coordinates to the sensor frame of reference. $K$ is the intrinsic matrix comprising of the focal lengths $f_{x}$ and $f_{y}$, and principal point $[u_{0},v_{0}]$, obtained through sensor calibration. In addition to the intrinsic matrix $K$, lens distortion coefficients $\Psi$ also need to be computed when calibrating event cameras defined by
@@ III-B Unsupervised Clustering of Events
Intuitively, $\epsilon_{s}$ is not unique when tackling different sensor sizes and calibration pattern variants of different geometric properties (i.e. height, width, and diagonal spacing). To circumvent this limitation, the spatial and temporal stamps of $e_{k}\in\mathbf{E}_{k}$ are normalized respect to the sensor resolution ($H\times W$) and the spatiotemporal window time step size $\Delta_{k}$ as $\hat{e}_{k}=(\frac{x}{W}\text{ }\frac{y}{H},\frac{t_{k}-t_{1}}{\Delta_{k}})$, where $\hat{e}_{k}$ are the normalized events. Consequently, the input to ST-DBSCAN are the normalized events. Normalizing the events’ spatial and temporal stamps makes the clustering resolution agnostic, independent of the accumulation time step size, and eventually a fixed $\epsilon_{s}$ is utilized for different sensor models. Fig. 4 shows the clustering performance on $\mathbf{E}_{K}$ and $\hat{\mathbf{E}}_{K}$ both comprising of 4000 elements for two sensor models of different resolutions, DAVIS346 and DVXplorer. Note that the same value of $\epsilon_{s}$ is identical for $\hat{\mathbf{E}}_{k}$ as the events are unified under the same domain range for both sensors while ST-DBSCAN fails on raw events fails of DVXplorer.
The output of ST-DBSCAN are the sets of clustered events $C=\{{I_{1},I_{2},...I_{J}}\}$, with $\mathbf{\hat{E}}_{K}$ partitioned as $\mathbf{\hat{E}}_{K}=I_{1}\cup I_{2}\cup...\cup I_{J}$, where $I_{j}\in C$ and $J$ is the total number of clusters. In our approach, we have strictly utilized at least 4000 events to cluster $\mathbf{\hat{E}}_{K}$ with a time step of 33 ms between the spatiotemporal windows. We avoided forming spatiotemporal windows based on fixed time intervals since at certain instances a very small relative motion is encountered and no sufficient features can be extracted from $\hat{e}_{k}$. On the other hand, $\epsilon_{s}$ and $\epsilon_{t}$ are set to 0.015 providing optimal clustering performance. It is also important to highlight that the event polarities are ignored in the clustering algorithm.
@@ IV Feature Extraction and Calibration
Figure 6: The body frame $\mathcal{F}_{bj}$ defined for each cluster in the spatiotemporal window with the cylinder centroid defined by $\mathbf{t}_{h,k,t}$.
@@ IV-A Efficient Reweighted Least Squares
8:         Find $R_{\beta,\alpha}$ and $\mathbf{t}_{u_{j},v_{j},t_{ref}}$
@@
13:         Check for convergence
14:     end while
15:end for
After obtaining the sets of clustered events $C$, each of the clustered sets $\{I_{1},I_{2},...,I_{J}\}$ are fed to eRWLS to efficiently fit and extract the centroid of the dense cylinders formed by the observed circles grid. We first highlight the motivation of eRWLS by showing the remaining presence of outliers, where the cluster centers drift away from the true cylinder centroid, see Fig. 5. Accordingly, eRWLS attempts to predict the desired centroids even in the presence of persisting sensor noise. Algorithm 1 defines the algorithmic procedure of eRWLS for estimating the cylinder centroids. It is also worth mentioning that the predicted centroids by eRWLS are not the actual image points to be considered for the calibration optimization. Instead, the estimated centroids are fed to the modified hierarchical clustering algorithm for detecting the circles grid before the calibration optimization.
As depicted in Fig. 3, the events for the calibration targets follow a cylinder rotated around the $x$ the $y$ axes of the spatiotemporal window $\mathcal{F}_{S}$ with angles $\beta_{j}$ and $\alpha_{j}$, respectively. We define a body frame for each of cylinder $\mathcal{F}_{Bj}$ of basis $\bm{[b_{x},b_{y},b_{t}]}$, where $\bm{b_{t}}$ resembles a monotonically increasing time dimension, see Fig. 6. Accordingly, we define the transformation, ${}^{B}_{S}{\mathcal{T}}_{j}=[R_{\beta,\alpha}|\mathbf{t}_{u_{j},v_{j}}]$ transforming the events spatial and temporal stamps from $\mathcal{F}_{S}$ to $\mathcal{F}_{Bj}$ as
${}^{j}\tilde{e}_{k}=\begin{bmatrix}\tilde{x}_{k}\\ \tilde{y}_{k}\\ \tilde{t}_{k}\end{bmatrix}=\underbrace{\begin{bmatrix}R_{\beta_{j},\alpha_{j}}\text{ }|\text{ }\mathbf{t}_{u_{j},v_{j},t_{ref}}\end{bmatrix}}_{{}^{B}_{S}{\mathcal{T}}_{j}}\begin{bmatrix}x_{k}\\ y_{k}\\ t_{k}\\ 1\end{bmatrix},$ (4)
where ${}^{j}\tilde{e}_{k}$ are the events defined in $\mathcal{F}_{Bj}$, $R_{\beta_{j},\alpha_{j}}$ is the cylinder orientation, and $\mathbf{t}_{u_{j},v_{j},t_{ref}}$ is the translation vector pointing from $\mathcal{F}_{S}$ origin to the cylinder centroid. Note that the cylinder height is constrained at reference time $t_{ref}=t_{1}$ and is not optimized. Meaning, all cylinder centroids are set to their caps at $t_{1}$, see Fig. 7. This is of paramount importance to ensure that all cylinder centroids are synchronized in time. To optimize the other the cylinder parameters, $\bm{\Omega_{j}}=[r_{j},\text{ }u_{j},\text{ }v_{j},\text{ }\beta_{j},\text{ }\alpha_{j}]$, a nonlinear least square problem can be constructed to minimize the squared residuals $\bm{\xi}$ formulated as
Figure 7: Extracted cylinder centroids by eRWLS. Notice that all centroids are synchronized at $t_{ref}=t_{1}$.
@@
where $u_{j}$ and $v_{j}$ are the cylinder centroid, while $r_{j}$ is the cylinder radius. However, the noise in event stream as shown in Fig. 5, will introduce outliers to the optimization framework which can significantly affect the optimization results. A method that provides solutions to the least squares problem that is robust against noise is the weighted least squares, but the weights for each element in the residual vector need to be provided in prior. On the other hand, iteratively reweighted least squares (IRLS) [37] can iteratively recompute the weight vector but suffers from convergence difficulties and leads to undesired computational complexities. To alleviate these restrictions, we show that the probability density function (PDF) of $\bm{\xi}$ is dominated with a Gaussian distribution and the weights $\mathbf{w}$ can be implicitly updated by a normally distributed vector as
$\mathbf{w}_{j}=\dfrac{1}{\sigma\sqrt{2\pi}}e^{-(\bm{\xi}\mu)^{2}/{2\sigma^{2}}},$ (6)
where $\mu$ and $\sigma$ are the mean and standard deviation of $\bm{\zeta}$. Accordingly, the weights for the outliers are severely weakened and the cost function in Eq. 5 is modified to
@@
One important aspect of the optimization process is the initial condition $\omega_{j}$ of $\Omega_{j}$. To maintain stability and fast convergence, $\omega_{j}$ is parameterized by the median of the transformed spatial stamps ${}^{j}\tilde{e}_{k}$ along with their norm $\bar{r}_{j}$ as $\bar{u}_{j}$ and $\bar{v}_{j}$, while $\beta_{j}$ and $\alpha_{j}$ are initialized as zeros. The optimization is run with a variable differential step computed from the jacobian $\mathbf{J}$ [38], which has been found analytically using MATLAB symbolic toolbox [39].

### v2

# This paper has been accepted for publication at IEEE Transactions on Image Processing (TIP), 2024. ©IEEE

@@ Abstract
Event cameras triggered a paradigm shift in the computer vision community delineated by their asynchronous nature, low latency, and high dynamic range. Calibration of event cameras is always essential to account for the sensor intrinsic parameters and for 3D perception. However, conventional image-based calibration techniques are not applicable due to the asynchronous, binary output of the sensor. The current standard for calibrating event cameras relies on either blinking patterns or event-based image reconstruction algorithms. These approaches are difficult to deploy in factory settings and are affected by noise and artifacts degrading the calibration performance. To bridge these limitations, we present E-Calib, a novel, fast, robust, and accurate calibration toolbox for event cameras utilizing the asymmetric circle grid, for its robustness to out-of-focus scenes. E-Calib introduces an efficient reweighted least squares (eRWLS) method for feature extraction of the calibration pattern circles with sub-pixel accuracy and robustness to noise. In addition, a modified hierarchical clustering algorithm is devised to detect the calibration grid apart from the background clutter. The proposed method is tested in a variety of rigorous experiments for different event camera models, on circle grids with different geometric properties, on varying calibration trajectories and speeds, and under challenging illumination conditions. The results show that our approach outperforms the state-of-the-art in detection success rate, reprojection error, and pose estimation accuracy.
@@ I Introduction
Over the past decades, imaging sensors have evolved at a rapid pace providing solutions for various intelligent perception algorithms. The event camera is a bio-inspired sensor whose pixels asynchronously fire events upon change of light intensity. Each generated event records the coordinates of the pixel spiking, the time at which the event occurred, and the sign of the intensity change. The output of an event camera is therefore not frames at constant time intervals but rather an asynchronous stream of events in space and time. Due its asynchronous and differential nature, an event camera has three key advantages for robot manufacturing: sub-millisecond latency, microsecond resolution, and very high dynamic range [1, 2, 3]. These features grant the sensor robustness against motion blur and poorly lighted environments, marking a new paradigm for vision sensors. With such unprecedented capabilities in hand, event-based vision unveiled its impact in space robotics [4, 5, 6], robotic manufacturing [7], fast object detection and tracking [8, 9, 10, 11], autonomous navigation [12, 13, 14], and indoor positioning systems [15].
@@
More sophisticated event camera models, also called Dynamic Active-Pixel Vision Sensor (DAVIS) [18], are accompanied by frame-based imaging capabilities, where the active-pixel sensor (APS) frames are utilized for calibration using conventional calibration techniques. As event-based vision is gradually gaining maturity, recent models of event cameras are rather manufactured as stand-alone sensors. Thus, Huang et al. [19] proposed a multi-segment optimization method to obtain the event camera intrinsics from the raw events using the asymmetric circle grid for its robustness against defocused scenes. However, the approach utilizes the fired raw events for the calibration optimization. Hence, sub-pixel localization accuracy of the calibration feature points is not attained degrading the calibration accuracy. This ultimately instigates a surging demand for a robust, event-driven calibration tool to obtain the intrinsic parameters of event cameras.
@@
    We validated the proposed work in rigorous experiments on two event camera models, using different calibration patterns, under challenging lighting conditions, and on three calibration trajectories at different speeds. The results show that our approach outperforms the state-of-the-art in terms of detection success rate, reprojection error, extrinsics estimation accuracy, and computational efficiency.
@@ I-A Related Work
Although image processing techniques have advanced in the past decades, the unconventional nature of event cameras inhibits the applicability of such algorithms for event-based vision. Still, knowledge of the sensor intrinsic parameters remains necessary for developing robust event-based perception algorithms. The first method developed for calibrating event cameras utilized blinking calibration patterns [16]. Since event-based vision sensors require light intensity variation to generate events, blinking screens alleviate this restriction by consistently firing events resembling checkerboard and circle calibration grids. Nonetheless, such approach needs unnecessary instrumentation as specialized screens are required to render the calibration pattern at specific frequencies [17]. More importantly, it is usually desired to move the calibration pattern instead of the event camera during calibration. This is the case when the imaging sensor is mounted on unmanned aerial and ground vehicles [29], where manipulating the blinking screen is not feasible. Due to the limitations of blinking calibration patterns, recent advances in deep learning have been utilized to reconstruct grayscale frames, where conventional calibration methods are utilized to obtain the event camera intrinsics from the reconstructed frames. The first deep neural network to provide such utility is the E2VID network [30] where E2Calib [17] was proposed utilizing E2VID for the purpose of sensor calibration. Consequently, several improved networks have been developed: 1- Spade-E2VID by Cadena et al. [31] and 2- FireNet by Scheerlinck et al. [32]. However, the reconstructed images can suffer from artifacts as the sensor sweeps the calibration pattern degrading the feature extraction and calibration performance. This also shows that traditional image-based perception algorithms do not maximize the potential of event-based vision. Since traditional image processing techniques for event camera calibration do not provide the sufficient calibration accuracy, Huang et al. [19] proposed a multi-segment based optimization to obtain the sensor intrinsics directly from events using the asymmetric circle grid for its robustness against defocused scenes [33]. The approach associates events to their respective circles using density-based spatial clustering (DBSCAN) and introduces soft and hard feature extraction methods to differentiate events generated by the calibration pattern circles from background noise. The incorporated feature extraction methods rely on circle fitting for each cluster to pass events along the fitted circles to a multi-segment optimization problem to obtain the intrinsics. The drawback of such method is the presence of noisy events close to the observed circles, which affect the circle fitting accuracy leading to false detections and degraded calibration performance. In addition, the method utilizes the raw events as the control points for the calibration optimization instead of the circle centers of the circle targets. Hence, sub-pixel localization accuracy of the control points is not achieved with the edge events, which is directly correlated with degraded calibration accuracy.
Figure 1: A high-level block diagram of the proposed calibration method. First, the raw events $\mathbf{E}_{k}$ are acquired from the event camera and are clustered to their respective circles to the clustered event sets $C=\{{I_{1},I_{2},...I_{J}}\}$ using ST-DBSCAN. Second, the centers of the circles are robustly extracted by means of eRWLS and the pattern is distinguished from the background using modified hierarchical clustering. Finally, the calibration optimization is performed using the predicted circles 2D image points $\mathbf{U}_{M}$ to obtain the event camera intrinsics.
@@ II Framework Overview
An illustrative block diagram of our proposed calibration approach is shown in Fig. 1, involving a moving event camera observing an asymmetric circle grid. We chose the asymmetric circle grid as the calibration pattern for two reasons. First, the asymmetric circles calibration pattern is robust against out of focus scenes [33]. Second, the features generated from the circles are motion invariant, unlike checkerboards and AprilTag Markers where edges parallel to camera motion do not fire events [34], see Fig. 2. As the event camera sweeps the calibration pattern of known $M$ circles, the fired events need to be associated to their corresponding circles. Thus, we devise spatiotemporal density-based clustering with applications to noise (ST-DBSCAN) [35] to cluster the asynchronous events to their respective circular features, outlined in section III-B. Unlike DBSCAN [36] which ignores the timestamps of the events, ST-DBSCAN performs optimal event-based clustering due to its capability to operate on spatiotemporal data and adhere to the sparse output of the event camera. It is worth mentioning that ST-DBSCAN requires predefined parameters $\epsilon_{s}$ and $\epsilon_{t}$, defined as the neighborhood search radius, which require tuning for different event camera models. To circumvent this limitation, the asynchronous events are normalized with respect to the sensor resolution prior to ST-DBSCAN. This is essential to perform clustering in a unified spatiotemporal domain and ensure consistent clustering performance even for different sensor sizes.
@@
After successfully associating the events to their respective circles, the image centers of the calibration targets need to be extracted with sub-pixel localization accuracy to optimally calibrate the sensor. Such feature is not obtained by utilizing the raw events as control points for calibration optimization. A naive approach is to rely on ST-DBSCAN cluster centers as the feature points of the calibration targets. However, due to the inherent noisy nature of the event camera, the cluster centers drift away from the actual circle centers. Even though ST-DBSCAN is characterized with robustness to noise, noisy events can still be mapped to the circle targets, especially when they are close to the real events. On the other hand, the observed circular features form dense slanted cylinders in spatiotemporal space due to monotonically increasing event stamps, preventing the applicability of image-based feature extraction methods. To extract the image points of the pattern circles, each clustered set of events is first fed to the efficient weighted least squares (eRWLS) algorithm, discussed in section IV-A, where cylinders are robustly fitted to the events clusters while penalizing the noisy events close to the observed circles. It is also essential to highlight that eRWLS assumes the probability density function (PDF) of the residual vector is dominated by a normal distribution. Accordingly, eRWLS efficiently reweights the residual vector while penalizing the outlier weights iteratively until convergence.
Finally, since background clutter can be observed, a modified hierarchical clustering algorithm, detailed in section IV-B, is devised to extract the calibration pattern in the image plane apart from the background given the geometric properties of the asymmetric circle grid. Note that we employed the hierarchical clustering algorithm described in [37]. However, the algorithm is not directly applicable since hierarichal clustering on the detected circle centers fails since they do not reflect the geometric representation of the calibration pattern. Thus, we transform the input data to conform to the uniform geometric representation of the circles in the calibration pattern. This adjustment ensures clear differentiation from background clutter.
As soon as the image centers are identified, they are fed to the calibration optimizer along with the pattern geometric properties to obtain the sensor intrinsic parameters.
@@ III-A Event-Based Vision
$s\begin{bmatrix}u\\ v\\ 1\end{bmatrix}=K\underbrace{\begin{bmatrix}R\text{ }T\end{bmatrix}}_{{}^{P}_{C}\mathcal{T}}\begin{bmatrix}x\\ y\\ z\\ 1\end{bmatrix},$ (2)
where $x$, $y$, $z$ represent the world coordinates of a point in 3D space mapped to image coordinates $u$ and $v$. ${}^{P}_{C}\mathcal{T}$ is the transformation from world, i.e. calibration pattern, coordinates to the event camera frame of reference defined by the rotation matrix $R$ and translation vector $T$. $K$ is the intrinsic matrix comprising of the focal lengths $f_{x}$ and $f_{y}$, and principal point $[u_{0},v_{0}]$, obtained through sensor calibration. In addition to the intrinsic matrix $K$, lens distortion coefficients $\Psi$ also need to be computed when calibrating event cameras defined by
@@ III-B Unsupervised Clustering of Events
Intuitively, $\epsilon_{s}$ is not unique when tackling different sensor sizes and calibration pattern variants of different geometric properties (i.e. height, width, and diagonal spacing). To circumvent this limitation, the spatial and temporal stamps of $e_{k}\in\mathbf{E}_{k}$ are normalized with respect to the sensor resolution ($H\times W$) and the spatiotemporal window time step size $\Delta_{k}$ as $\hat{e}_{k}=(\frac{x}{W}\text{ }\frac{y}{H},\frac{t_{k}-t_{1}}{\Delta_{k}})$, where $\hat{e}_{k}$ are the normalized events. Consequently, the input to ST-DBSCAN are the normalized events. Normalizing the events’ spatial and timestamps makes the clustering resolution agnostic, independent of the accumulation time step size, and eventually a fixed $\epsilon_{s}$ is utilized for different sensor models. Fig. 4 shows the clustering performance on $\mathbf{E}_{K}$ and $\hat{\mathbf{E}}_{K}$ both comprising of 4000 events for two sensor models of different resolutions, DAVIS346 and DVXplorer. Note that the same value of $\epsilon_{s}$ is identical for $\hat{\mathbf{E}}_{k}$ as the events are unified under the same domain range for both sensors while ST-DBSCAN fails on the DVXplorer raw events.
The output of ST-DBSCAN are the sets of clustered events $C=\{{I_{1},I_{2},...I_{J}}\}$, with $\mathbf{\hat{E}}_{K}$ partitioned as $\mathbf{\hat{E}}_{K}=I_{1}\cup I_{2}\cup...\cup I_{J}$, where $I_{j}\in C$ and $J$ are the total number of clusters. Note that the spatial coordinates are in the original image plane and not the normalized coordinates. This is essential to extract the circle centers in the image plane for calibration and not in normalized coordinates. In addition, the total number of events utilized to formulate a spatiotemporal window of clustering is a hyperparameter. We avoided forming spatiotemporal windows based on fixed time intervals since at certain instances a very small relative motion is encountered and no sufficient features can be extracted from $\hat{e}_{k}$. In addition, large number of events can lead to curved cylinders generated from the observed circles. This tends to have insignificant effect on the calibration accuracy due to the weighing factor placed in the eRWLS cost function. Yet, the number of events is considered a hyperparameter to ensure the highest calibration accuracy can be obtained. In our experiments, we utilized at least 4000 events to cluster $\mathbf{\hat{E}}_{K}$ with a time step of 33 ms between the spatiotemporal windows during experiments. This was sufficient to obtain the required calibration accuracy. On the other hand, $\epsilon_{s}$ and $\epsilon_{t}$ are set to 0.015 providing optimal clustering performance. It is also important to highlight that the event polarities are ignored in the clustering algorithm.
@@ IV Feature Extraction and Calibration
Figure 6: The body frame $\mathcal{F}_{bj}$ defined for each cluster in the spatiotemporal window with the cylinder center defined by $\mathbf{t}_{b}$.
@@ IV-A Efficient Reweighted Least Squares
8:         Find $R_{\beta,\alpha}$ and $\mathbf{t}_{b}$
@@
13:     end while
14:end for
After obtaining the sets of clustered events $C$, each of the clustered sets $\{I_{1},I_{2},...,I_{J}\}$ are fed to eRWLS to efficiently fit and extract the center of the dense cylinders formed by the observed circles grid. We first highlight the motivation of eRWLS by showing the remaining presence of outliers, where the cluster centers drift away from the true cylinder center, see Fig. 5. Similarly, utilizing ordinary least squares (OLS) tends to be affected by the presence of noise leading to degraded detection performance of the circle centers. In contrast, eRWLS attempts to predict the desired centers even in the presence of persisting sensor noise. Algorithm 1 defines the algorithmic procedure of eRWLS for estimating the cylinder centers. It is also worth mentioning that the predicted centers by eRWLS are not the actual image points to be considered for the calibration optimization. Instead, the estimated centers are fed to the modified hierarchical clustering algorithm for detecting the circles grid before the calibration optimization.
As depicted in Fig. 3, the events for the calibration targets follow a cylinder rotated around the $x$ the $y$ axes of the spatiotemporal window $\mathcal{F}_{S}$ with angles $\beta_{j}$ and $\alpha_{j}$, respectively. We define a body frame for each of cylinder $\mathcal{F}_{Bj}$ of basis $\bm{[b_{x},b_{y},b_{t}]}$, where $\bm{b_{t}}$ resembles a monotonically increasing time dimension, see Fig. 6. Accordingly, we define the transformation, ${}^{B}_{S}{\mathcal{T}}_{j}$ described by the rotation $R_{\beta_{j},\alpha_{j}}$ and $\mathbf{t}_{b}=[u_{j},v_{j},t_{ref}]$ transforming the events coordinates from $\mathcal{F}_{S}$ to $\mathcal{F}_{Bj}$ as
${}^{j}\tilde{e}_{k}=\begin{bmatrix}\tilde{x}_{k}\\ \tilde{y}_{k}\\ \tilde{t}_{k}\end{bmatrix}=\underbrace{\begin{bmatrix}R_{\beta_{j},\alpha_{j}}&\mathbf{t}_{b}\\ 0&1\end{bmatrix}}_{{}^{B}_{S}{\mathcal{T}}_{j}}\begin{bmatrix}x_{k}\\ y_{k}\\ t_{k}\\ 1\end{bmatrix},$ (4)
where ${}^{j}\tilde{e}_{k}$ are the events defined in $\mathcal{F}_{Bj}$, $R_{\beta_{j},\alpha_{j}}$ is the cylinder’s orientation, and $\mathbf{t}_{b}$ is the translation vector pointing from $\mathcal{F}_{S}$ origin to the cylinder center. Note that the cylinder height is constrained at reference time $t_{ref}=t_{1}$ and is not optimized. Meaning, all cylinder centers are set to their caps at $t_{1}$, see Fig. 7. This is of paramount importance to ensure that all cylinder centers are synchronized in time. Note that any reference time can be used ranging from $t_{1}$ to $t_{k}$ as long as all the fitted circles are synchronized to the chosen $t_{ref}$.
To optimize the other the cylinder parameters, $\bm{\Omega_{j}}=[r_{j},\text{ }u_{j},\text{ }v_{j},\text{ }\beta_{j},\text{ }\alpha_{j}]$, a nonlinear least square problem can be constructed to minimize the squared residuals $\bm{\xi}$ formulated as
Figure 7: Extracted cylinder centers by eRWLS. Notice that all centers are synchronized at $t_{ref}=t_{1}$.
@@
where $u_{j}$ and $v_{j}$ are the pixel coordinates of the cylinder’s center, while $r_{j}$ is the cylinder radius. However, the noise in event stream as shown in Fig. 5, will introduce outliers to the optimization framework which can significantly affect the optimization results. A method that provides solutions to the least squares problem that is robust against noise is the weighted least squares, but the weights for each element in the residual vector need to be provided in prior. On the other hand, iteratively reweighted least squares (IRLS) [38] can iteratively recompute the weight vector but suffers from convergence difficulties and leads to undesired computational complexities. To alleviate these restrictions, we show that the probability density function (PDF) of $\bm{\xi}$ is dominated with a Gaussian distribution and the weights $\mathbf{w}$ can be implicitly updated by a normally distributed vector as
$\mathbf{w}_{j}=\dfrac{1}{\sigma\sqrt{2\pi}}\exp{(-\frac{\bm{\xi}\mu^{2}}{{2\sigma^{2}}})},$ (6)
where $\mu$ and $\sigma$ are the mean and standard deviation of $\bm{\xi}$. Accordingly, the weights for the outliers are severely weakened and the cost function in Eq. 5 is modified to
@@
One important aspect of the optimization process is the initial condition $\omega_{j}$ of $\Omega_{j}$. To maintain stability and fast convergence, we define $\omega_{j}$ to be the median of the transformed spatial stamps ${}^{j}\tilde{e}_{k}$ along with their norm $\bar{r}_{j}$ as $\bar{u}_{j}$ and $\bar{v}_{j}$, while $\beta_{j}$ and $\alpha_{j}$ are initialized as zeros. The optimization is run with a variable differential step computed from the jacobian $\mathbf{J}$ [39], which has been found analytically using MATLAB symbolic toolbox [40].